a) The commodity and money markets of Juja farm, a hypothetical economy are as given below: Commodity Market Y = C + I C = 1000 + 0.8Y I = 2000 – 0.75r Money Market Lt = 0.25Y (Transactions and precautionary demand for money function) Ls = 1000 – 0.5r (Speculative demand for money function) Ms = 3200 (Money supply function) Required: i) Mathematically derive both IS and LM curves. (6 marks) ii) Derive the equilibrium level of Income and rate of interest. (4 marks) iii) If the money supply is increased by 80, what would be the effect on the equilibrium level of income and rate of interest in Juja farm economy? 


1
Expert's answer
2021-12-09T07:45:54-0500

i) IS and LM curves equations are:

IS: "Y = C + I = 1000 + 0.8Y I + 2000 \u2013 0.75r = 3000 + 0.8Y - 0.75r,"

0.2Y = 3000 - 0.75r,

Y = 15000 - 3.75r.

LM: Md = Ms,

"0.25Y + 1000 - 0.5r = 3200,"

Y = 8800 + 2r.

ii) The equilibrium level of Income and rate of interest are:

15000 - 3.75r = 8800 + 2r,

5.75r = 6200,

r = 1078.26,

"Y = 8800 + 2\u00d71078.26 = 10,956.52."

iii) If the money supply is increased by 80, then both the equilibrium level of income and rate of interest in Juja farm economy will increase.